Using High-Low to Calculate Predicted Total Variable Cost and Total Cost for Budgeted Output Pizza Vesuvio...

  1. Using High-Low to Calculate Predicted Total Variable Cost and Total Cost for Budgeted Output

    Pizza Vesuvio makes specialty pizzas. Data for the past 8 months were collected:

    Month Labor Cost Employee Hours
    January $6,900 410
    February $9,799 600
    March $8,040 680
    April $9,687 660
    May $8,390 530
    June $7,431 400
    July $9,390 620
    August $7,350 360

    Assume that this information was used to construct the following formula for monthly labor cost.

    Total Labor Cost = $6572 + ($2.16 x Employee Hours)

    Required:

    Assume that 705 employee hours are budgeted for the month of September. Use the total labor cost formula for the following calculations:

    1. Calculate total variable labor cost for September. Round your answer to nearest whole number.
    $

    2. Calculate total labor cost for September. Round your answer to nearest whole number.
    $
Answer #1

1) Total variable labor cost = 2.16*705 = 1523

2) Total labor cost = Variable cost+Fixed cost = 1523+6572 = 8095